# Learning Internal Biological Neuron Parameters and Complexity-Based Encoding for Improved Spiking Neural Networks Performance

### v1

@@ Abstract
This study introduces a novel approach by replacing the traditional perceptron neuron model with a biologically inspired probabilistic meta neuron, where the internal neuron parameters are jointly learned, leading to improved classification accuracy of spiking neural networks (SNNs). To validate this innovation, we implement and compare two SNN architectures: one based on standard leaky integrate-and-fire (LIF) neurons and another utilizing the proposed probabilistic meta neuron model. As a second key contribution, we present a new biologically inspired classification framework that uniquely integrates SNNs with Lempel-Ziv complexity (LZC) a measure closely related to entropy rate. By combining the temporal precision and biological plausibility of SNNs with the capacity of LZC to capture structural regularity, the proposed approach enables efficient and interpretable classification of spatiotemporal neural data, an aspect not addressed in existing works. We consider learning algorithms such as backpropagation, spike-timing-dependent plasticity (STDP), and the Tempotron learning rule. To explore neural dynamics, we use Poisson processes to model neuronal spike trains, a well-established method for simulating the stochastic firing behavior of biological neurons. Our results reveal that depending on the training method, the classifier’s efficiency can improve by up to 11.00%, highlighting the advantage of learning additional neuron parameters beyond the traditional focus on weighted inputs alone.
keywords:
Spiking Neural Networks , Learning Algorithms , Neuron Models , Lempel-Ziv Complexity , Spike Trains, Classification
journal: Title
{highlights}
Joint neuron parameter and synaptic learning improves classification accuracy
SNN-LZC framework enables efficient, interpretable spatiotemporal classification
Validated across learning rules, the method shows robustness and flexibility
Supports biologically grounded models and rules for biosignal classification
Findings inform biologically inspired SNN design for classifications tasks
@@ 1 Introduction
Spiking Neural Networks (SNNs) have emerged as promising candidates for addressing the limitations of conventional artificial neural networks (ANNs), particularly in real-world applications where temporal precision and energy efficiency are critical [1]. Unlike traditional ANNs, SNNs more closely approximate the behavior of biological neurons by transmitting information through discrete spike events, thus capturing both the spatial and temporal dynamics of neural activity [3, 2]. This biologically grounded architecture positions SNNs as potential computational models capable of approaching the efficiency and adaptability of the human brain [4].
Despite their promise, training SNNs remains a significant challenge due to their non-linear, event-driven dynamics and the non-differentiable nature of spike generation [6, 7]. Traditional gradient-based methods such as backpropagation, which iteratively minimize a loss function by computing gradients of continuous activation values, are not well-suited to the discrete and temporally sparse nature of spiking activity [5]. Moreover, these conventional arithmetic-based learning rules are often biologically implausible and inefficient when applied to spiking architectures [8]. Instead, the optimization of intrinsic neuron parameters and the adoption of biologically inspired temporal coding schemes play a critical role in improving the performance and interpretability of SNNs. Temporal coding, in particular, has been shown to transmit information more efficiently than rate coding, potentially reducing the number of spikes required for accurate representation and decision-making [9].
@@
In this study, we propose a novel modification to the spiking neural network by replacing the traditional perceptron neuron model with a biologically inspired meta neuron, in which internal parameters are jointly learned during training. We hypothesize and demonstrate that this biologically motivated enhancement can significantly improve SNN classification accuracy. To validate this, we implement two SNN architectures: one based on standard leaky integrate-and-fire (LIF) neurons [15] and the other on the proposed meta neuron model [16]. As a second key contribution, we introduce a novel biologically inspired classification framework that integrates SNNs with Lempel-Ziv complexity. By leveraging the temporal precision and biological plausibility of SNNs together with LZC’s ability to quantify structural complexity, this framework provides efficient and interpretable analysis of spatiotemporal neural patterns for classification tasks, namely a capability not present in prior approaches. We further evaluate multiple learning algorithms, including backpropagation [5], spike-timing-dependent plasticity (STDP) [17], and the Tempotron learning rule [18, 19].
@@
$\textbf{x}=[x_{1},x_{2},\cdots,x_{n}]\in\mathbb{R}^{n}$ synaptic weight vector
$V_{m}(t)\in\mathbb{R}$ membrane potential at time $t$
$\tau_{m}\in\mathbb{R}^{+}$ membrane time constant
$I(t)\in\mathbb{R}$ input current at time $t$
$\tau_{m}\in\mathbb{R}^{+}$ membrane time constant
$t_{f}\in\mathbb{R}^{+}$ firing time of neuron
$S_{i}(t)\in\mathbb{Z}$ binary-valued spike train (0 or 1 at any given time)
$t_{i}^{k}\in\mathbb{R}^{+}$ spike times (real-valued, positive)
$\eta\in\mathbb{R}^{+}$ general learning rate (positive real number) for weight update in supervised learning and gradient-based optimization
$\eta_{+},\eta_{-}\in\mathbb{R}^{+}$ STDP time constant (positive real)
$A_{-},A_{+}\in\mathbb{R}^{+}$ amplitude of STDP weight changes (positive real)
$t_{pre},t_{post}\in\mathbb{R}^{+}$ pre- and postsynaptic spike time (positive real)
$t_{spike}\in\mathbb{R}^{+}$ spike timing
$S_{pre}(t),S_{post}(t)\in{0,1}$ binary spike trains
$K(t)\in\mathbb{R}\longrightarrow\mathbb{R}$ kernel function, maps time to real value
$t_{i}^{target},t_{i}^{actual}\in\mathbb{R}^{+}$ targeted and ctual spike times
$\mathcal{E}\in\mathbb{R}^{+}$ error function (positive real)
$\frac{\partial E}{\partial w_{i}}\in\mathbb{R}$ gradient of error
Table 1: Basic notation used in the study.
@@ 2 Notation
Figure 1: Comparison of the standard Leaky Integrate-and-Fire (LIF) neuron model and the extended meta-neuron architecture. The LIF neuron integrates the input current $I(t)$ over time according to the leaky membrane equation $\tau\frac{dV}{dt}=-V+RI(t)$, where $\tau$ is the membrane time constant and $R$ is the membrane resistance. A spike is emitted when the membrane potential $V(t)$ exceeds a fixed threshold $V_{\text{th}}$, followed by a reset to $V_{\text{reset}}$. The meta-neuron preserves this core structure but incorporates dynamic, adaptive components. Specifically, the membrane time constant $\tau(t)$ and threshold $V_{\text{th}}(t)$ can vary over time, enabling richer and more flexible temporal dynamics. More generally, the meta-neuron can be interpreted as computing a parameterized function $y=f(\mathbf{x};\boldsymbol{\theta})$, where the input vector $\mathbf{x}$ is mapped to an output $y$ through parameters $\boldsymbol{\theta}$, which may be adapted online through feedback mechanisms or learning algorithms. This structure supports meta-learning and context-aware adaptation while maintaining biologically inspired spiking behavior.
@@ 6 Results
In our study, we introduce a novel spiking neural network framework that, to the best of our knowledge, is the first to integrate Lempel–Ziv Complexity with biologically inspired neuron models. We employ two distinct SNN architectures: one based on the widely used leaky integrate-and-fire neurons [15], and the other utilizing a biologically motivated meta neuron model [21]. This integration of LZC with SNNs provides a new classification approach, leveraging the temporal coding capabilities and neural plausibility of SNNs along with the quantification of structural complexity provided by LZC. The result is a robust and interpretable method for analyzing spatio-temporal data, an aspect that remains largely unexplored in the current SNN literature. To optimize performance and evaluate learning dynamics, we implement and compare several training algorithms, including backpropagation-based methods [5], spike-timing-dependent plasticity [17], and the Tempotron learning rule [18, 19].
@@
Table 2 presents the classification accuracy results for spiking neural networks based on the leaky integrate-and-fire neuron model, using inputs generated as Poisson spike trains. Our findings indicate that optimizing learning and neuronal parameters can significantly improve classification accuracy. In particular, the application of Backpropagation and spike-timing-dependent plasticity led to accuracy improvements of up to 11.00% when both neuron parameters and the learning rate were optimized, as compared to the baseline configuration where only the learning rate was trained. In contrast, the Tempotron learning rule exhibited a more modest improvement of approximately 4.00%. However, its baseline performance, achieved by training only the learning rate, was already high, reaching 95.00% accuracy.
Table 2: Accuracy results for SNN based on LIF model.
Learning Parameters Backpropagation Tempotron STPD
Number of neurons
Learning rate
85.50% 95.00% 88.50%
Number of neurons
Learning rate
Thresholds
Decays
96.50% (11.00% $\uparrow$) 99.00% (4.00% $\uparrow$) 99.50%(11.00% $\uparrow$)
@@
Learning Parameters Backpropagation Tempotron STPD
Number of neurons
Learning rate
79.00% 85.00% 87.00%
Number of neurons
Learning rate
Thresholds
Decays
90.00% (11.00% $\uparrow$) 92.00% (7.00% $\uparrow$) 91.50%(4.50% $\uparrow$)
In Table 3 the classification accuracy results for spiking neural networks based on meta-neuron is shown. Similar to the LIF-based architecture, it was observed that jointly optimizing both the neuron parameters and the learning rate led to a noticeable improvement in classification accuracy. Among the evaluated learning algorithms, Backpropagation yielded the most substantial gain, with an increase of up to 11.00% relative to the baseline configuration where only the learning rate was adjusted.
@@
Backpropagation and STDP benefited the most from intrinsic parameter optimization, achieving up to 11% accuracy gain relative to the baseline. Tempotron, while already showing a strong baseline performance, exhibited smaller improvements (up to 4–7%), indicating that its performance relies primarily on synaptic weight dynamics rather than intrinsic parameter tuning. The smaller improvement observed for the Tempotron can be explained by the nature of its learning rule, which relies primarily on adjusting synaptic weights to shape the decision boundary. As a result, parameter tuning (thresholds and decay constants) contributes less to performance compared to Backpropagation or STDP. Thus,the optimization configurations and their interpretation was summarized in Table 4.

### v2

# arXiv:2508.11674v2[cs.NE] 02 Mar 2026

@@ Abstract
This study introduces a novel learning paradigm for spiking neural networks (SNNs) by replacing the traditional perceptron-inspired neuron abstraction with biologically motivated neuron models in which not only synaptic weights but also intrinsic neuronal parameters are jointly optimized during training. We evaluate two SNN architectures, one based on standard leaky integrate-and-fire (LIF) neurons and another employing a meta-neuron mode, and compare settings with fixed versus learnable intrinsic parameters. As a second contribution, we propose a biologically inspired classification framework that integrates SNN dynamics with Lempel-Ziv complexity (LZC), a measure closely related to entropy rate, enabling efficient and interpretable classification of spatiotemporal spike data.
We consider multiple learning algorithms, including backpropagation with surrogate gradients, spike-timing-dependent plasticity (STDP), and the Tempotron rule, and train on spike trains generated from Poisson processes, which are widely adopted in computational neuroscience as a standard stochastic model for neuronal spike generation due to their analytical tractability and empirical relevance to irregular cortical firing patterns. Experiments show that learning intrinsic neuron parameters in addition to synaptic weights improves classification accuracy by up to 13.50 percentage points for LIF-based SNNs and up to 8.50 percentage points for meta-neuron models, relative to baselines where only network size and learning rate are tuned. Moreover, the proposed SNN-LZC classifier reaches accuracies up to 99.50% while maintaining sub-millisecond inference latency and competitive energy consumption.
In addition to empirical evaluation, we provide a quantitative theoretical justification for the observed gains by formalizing the effect of enlarging the hypothesis class through learnable intrinsic dynamics and deriving descent guarantees for intrinsic-parameter updates under standard smoothness (Lipschitz-gradient) assumptions. This analysis links neuron-intrinsic optimization to provable improvements in the surrogate objective and supports the practical relevance of jointly optimizing internal neuron dynamics and complexity-based readout.
keywords
Spiking Neural Networks, Learning Algorithms, Neuron Models, Lempel-Ziv Complexity, Spike Trains, Classification.
@@ 1 Introduction
Spiking Neural Networks (SNNs) have emerged as promising candidates to address the limitations of conventional artificial neural networks (ANNs), particularly in real-world applications where temporal precision and energy efficiency are critical [1]. Unlike traditional ANNs, SNNs more closely approximate the behavior of biological neurons by transmitting information through discrete spike events, thus capturing both the spatial and temporal dynamics of neural activity [2]. This biologically grounded architecture positions SNNs as potential computational models capable of approaching the efficiency and adaptability of the human brain [3].
Despite their promise, training SNNs remains a significant challenge due to their non-linear, event-driven dynamics and the non-differentiable nature of spike generation [5, 6]. Traditional gradient-based methods such as backpropagation, which iteratively minimize a loss function by computing gradients of continuous activation values, are not well-suited to the discrete and temporally sparse nature of spiking activity [4]. Moreover, these conventional arithmetic-based learning rules are often biologically implausible and inefficient when applied to spiking architectures [7, 8]. Instead, the optimization of intrinsic neuron parameters and the adoption of biologically inspired temporal coding schemes play a critical role in improving the performance and interpretability of SNNs [9]. Temporal coding, in particular, has been shown to transmit information more efficiently than rate coding, potentially reducing the number of spikes required for accurate representation and decision-making [10, 11].
@@
In this study, we propose a novel modification to the spiking neural network by replacing the traditional perceptron neuron model with a biologically inspired meta neuron, in which internal parameters are jointly learned during training. We hypothesize and demonstrate that this biologically motivated enhancement can significantly improve SNN classification accuracy. To validate this, we implement two SNN architectures: one based on standard leaky integrate-and-fire (LIF) neurons [17] and the other on the proposed meta neuron model [18, 19]. As a second key contribution, we introduce a novel biologically inspired classification framework that integrates SNNs with Lempel-Ziv complexity. By leveraging the temporal precision and biological plausibility of SNNs together with LZC’s ability to quantify structural complexity, this framework provides efficient and interpretable analysis of spatiotemporal neural patterns for classification tasks, namely a capability not present in prior approaches. We further evaluate multiple learning algorithms, including backpropagation [4], spike-timing-dependent plasticity (STDP) [20], and the Tempotron learning rule [21].
Table 1: Basic notation used in the study.
@@
$\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]\in\mathbb{R}^{n}$ Synaptic weight vector.
$V_{m}(t)\in\mathbb{R}$ Membrane potential at time $t$.
$\tau_{m}\in\mathbb{R}^{+}$ Membrane time constant.
$I(t)\in\mathbb{R}$ Input current at time $t$.
$t_{f}\in\mathbb{R}^{+}$ Neuron firing time.
$S_{i}(t)\in\{0,1\}$ Binary-valued spike train at time $t$.
$t_{i}^{k}\in\mathbb{R}^{+}$ Time of the $k$-th spike of neuron $i$.
$\eta\in\mathbb{R}^{+}$ Learning rate used for weight updates.
$\eta_{+},\eta_{-}\in\mathbb{R}^{+}$ STDP time constants.
$A_{+},A_{-}\in\mathbb{R}^{+}$ Amplitudes of STDP weight changes.
$t_{\mathrm{pre}},t_{\mathrm{post}}\in\mathbb{R}^{+}$ Pre- and postsynaptic spike times.
$K(t):\mathbb{R}\rightarrow\mathbb{R}$ Kernel function mapping time to a real value.
$t_{i}^{\mathrm{target}},t_{i}^{\mathrm{actual}}\in\mathbb{R}^{+}$ Target and actual spike times of neuron $i$.
$\mathcal{E}\in\mathbb{R}^{+}$ Error function value.
$\frac{\partial\mathcal{E}}{\partial w_{i}}\in\mathbb{R}$ Gradient of the error with respect to weight $w_{i}$.
@@ 2 Notation
[…]
Figure 1: Comparison of the standard Leaky Integrate-and-Fire (LIF) neuron model and the extended meta-neuron architecture. The LIF neuron integrates the input current $I(t)$ over time according to the leaky membrane equation $\tau\frac{dV}{dt}=-V+RI(t)$, where $\tau$ is the membrane time constant and $R$ is the membrane resistance. A spike is emitted when the membrane potential $V(t)$ exceeds a fixed threshold $V_{\mathrm{th}}$, followed by a reset to $V_{\mathrm{reset}}$. The meta-neuron preserves this core structure but incorporates dynamic, adaptive components. Specifically, the membrane time constant $\tau(t)$ and threshold $V_{\mathrm{th}}(t)$ can vary over time, enabling richer and more flexible temporal dynamics. More generally, the meta-neuron can be interpreted as computing a parameterized function $y=f(\mathbf{x};\boldsymbol{\theta})$, where the input vector $\mathbf{x}$ is mapped to an output $y$ through parameters $\boldsymbol{\theta}$, which may be adapted online through feedback mechanisms or learning algorithms.
@@ 6 Results
In our study, we introduce a novel spiking neural network framework that, to the best of our knowledge, is the first to integrate Lempel–Ziv Complexity with biologically inspired neuron models. We employ two distinct SNN architectures: one based on the widely used leaky integrate-and-fire neurons [17], and the other utilizing a biologically motivated meta neuron model [23]. This integration of LZC with SNNs provides a new classification approach, leveraging the temporal coding capabilities and neural plausibility of SNNs along with the quantification of structural complexity provided by LZC. The result is a robust and interpretable method for analyzing spatio-temporal data, an aspect that remains largely unexplored in the current SNN literature. To optimize performance and evaluate learning dynamics, we implement and compare several training algorithms, including backpropagation-based methods [4], spike-timing-dependent plasticity [20], and the Tempotron learning rule [21].
@@
Table 2 presents the classification accuracy results for spiking neural networks based on the leaky integrate-and-fire neuron model, using inputs generated as Poisson spike trains. Our findings indicate that optimizing both learning parameters and intrinsic neuron parameters can significantly improve classification accuracy. In particular, the application of backpropagation and spike-timing-dependent plasticity leads to gains from 82.50% to 96.00% (13.50 percentage points) and from 86.50% to 96.50% (10.00 percentage points), respectively, when thresholds and decay constants are learned in addition to the number of neurons and the learning rate. In contrast, the Tempotron learning rule exhibits a more modest improvement from 96.00% to 99.50% (3.50 percentage points). However, its baseline performance, achieved by training only the number of neurons and the learning rate, is already high, reaching 96.00% accuracy.
Table 2: Accuracy results for SNN based on LIF neuron model.
Learning algorithm Input Learning parameters Acc [%] P0 R0 F10 P1 R1 F11
BP Poisson Number of neurons Learning rate Thresholds Decays 96.00($13.50\uparrow$) 0.6809 0.9796 0.8033 0.9661 0.5588 0.7081
Number of neurons Learning rate 82.50 1.0000 0.0510 0.0971 0.5231 1.0000 0.6869
STDP Poisson Number of neurons Learning rate Thresholds Decays 96.50($10.00\uparrow$) 0.9333 1.0000 0.9655 1.0000 0.9314 0.9645
Number of neurons Learning rate 86.50 0.8515 0.8776 0.8643 0.8799 0.8529 0.8657
[…]
Number of neurons Learning rate 96.00 1.0000 0.0102 0.0202 0.5126 1.0000 0.6777
Notes. P0, R0, F10 denote Precision, Recall, and F1-score for Class 0, respectively. Analogously, P1, R1, F11 correspond to Precision, Recall, and F1-score for Class 1.
Table 3: Comparison of computational cost and time efficiency of the evaluated SNN based on LIF neuron model.
Learning algorithm Input Learning Rules $\overline{t}$ [ms] $t_{95}$ [ms] Throughput [samples/s] FLOPs fwd FLOPs bwd Total FLOPs Energy [J] Train FLOPs/s Eval FLOPs/s
BP Poisson Number of neurons Learning rate Thresholds Decays 0.3672 0.9249 2677.16 25600 25600 926 720 000 0.000927 1.36e+06 6.85e+07
Number of neurons Learning rate 0.5663 1.2282 1731.02 50176 50176 1 816 371 200 0.001816 1.35e+06 8.69e+07
STDP Poisson Number of neurons Learning rate Thresholds Decays 0.7678 1.2600 1279.03 65536 65536 1 192 755 200 0.001193 1.42e+07 8.38e+07
Number of neurons Learning rate 0.8998 1.1933 1102.03 1568 1568 56 761 600 0.000057 4.94e+05 1.73e+06
Tempotron Poisson Number of neurons Learning rate Thresholds Decays 0.1584 0.2333 6080.64 4096 4096 148 275 200 0.000148 3.32e+07 2.49e+07
Number of neurons Learning rate 0.1970 0.5060 4702.25 4096 4096 148 275 200 0.000148 3.48e+07 1.93e+07
Table 3 summarizes the computational cost and time efficiency of the LIF-based SNN for different learning rules and configurations. Across all methods, inference remains in the sub-millisecond regime, with average latency per sample ranging from $0.16$ ms to $0.90$ ms and 95th percentile latency below $1.3$ ms. This corresponds to a throughput between approximately $1.1\times 10^{3}$ and $6.1\times 10^{3}$ samples/s, indicating that all configurations are suitable for real-time processing. The extended configurations, in which intrinsic neuron parameters are learned in addition to synaptic weights, do not systematically increase the computational burden. For backpropagation, the extended model achieves a lower average latency ($0.37$ ms versus $0.57$ ms) and higher throughput (2677 samples/s vs. 1731 samples/s) despite using fewer FLOPs overall ( $9.27\times 10^{8}$ versus $1.82\times 10^{9}$). In contrast, the extended STDP configuration requires more compute per sample (up to 65 536 FLOPs in forward and backward passes) and a higher total FLOP budget than its baseline, but still maintains sub-millisecond latency. Tempotron-based LIF-SNNs achieve the highest throughput, with the extended variant reaching on average $0.16$ ms per sample and over 6000 samples/s, at a moderate total cost of $1.48\times 10^{8}$ FLOPs. Overall, the results show that substantial accuracy gains (Table 2) can be obtained without sacrificing real-time performance, and in some cases with improved speed–accuracy trade-offs.
In Table 4, the classification accuracy results for spiking neural networks based on the meta-neuron model are reported. Similar to the LIF-based architecture, jointly optimizing both intrinsic neuron parameters and the learning rate leads to a clear improvement in classification accuracy. Among the evaluated learning algorithms, backpropagation yields the largest gain, increasing accuracy from 89.00% in the baseline configuration (number of neurons and learning rate only) to 97.50% when thresholds and decay constants are also learned, i.e., by 8.50 percentage points. STDP shows a comparable trend, with an improvement from 88.50% to 95.50% (7.00 percentage points), while Tempotron achieves a more moderate gain from 89.50% to 93.00% (3.50 percentage ponits).
@@
Learning algorithm Input Learning parameters Acc [%] P0 R0 F10 P1 R1 F11
[…]
Number of neurons Learning rate 89.00 0.4900 1.0000 0.6577 0.0000 0.0000 0.0000
[…]
Number of neurons Learning rate 88.50 0.4900 1.0000 0.6577 0.0000 0.0000 0.0000
Tempotron Poisson Number of neurons Learning rate Thresholds Decays 93.00 (3.50 $\uparrow$) 0.9468 0.9082 0.9271 0.9151 0.9510 0.9327
Number of neurons Learning rate 89.50 0.7778 0.9286 0.8465 0.9157 0.7451 0.8216
Notes. P0, R0, F10 denote Precision, Recall, and F1-score for Class 0, respectively. Analogously, P1, R1, F11 correspond to Precision, Recall, and F1-score for Class 1.
Table 5: Comparison of computational cost and time efficiency of the evaluated SNN based on meta neuron model.
Learning algorithm Input Learning Rules $\overline{t}$ [ms] $t_{95}$ [ms] Throughput [samples/s] FLOPs fwd FLOPs bwd Total FLOPs Energy [J] Train FLOPs/s Eval FLOPs/s
BP Poisson Number of neurons Learning rate Thresholds Decays 0.9466 1.6320 1046.74 1800 1800 65 160 000 0.000065 5.80e+05 1.88e+06
Number of neurons Learning rate 0.7720 1.6226 1265.20 1568 1568 56 761 600 0.000057 6.51e+05 1.98e+06
STDP Poisson Number of neurons Learning rate Thresholds Decays 0.7720 1.6226 1265.20 1568 1568 56 761 600 0.000057 6.51e+05 1.98e+06
Number of neurons Learning rate 0.7720 1.6226 1265.20 1568 1568 56 761 600 0.000057 6.51e+05 1.98e+06
Tempotron Poisson Number of neurons Learning rate Thresholds Decays 0.6687 1.3717 1479.66 1568 1568 56 761 600 0.000057 4.28e+06 2.32e+06
Number of neurons Learning rate 0.5925 1.2652 1644.91 1568 1568 56 761 600 0.000057 4.43e+06 2.58e+06
[…]
Across all configurations, inference remains in the sub-millisecond regime, with average latency per sample ranging from 0.16 ms to 0.95 ms and throughput between approximately $1.1\times 10^{3}$ and $6.1\times 10^{3}$ samples/s. While the extended configurations with learned intrinsic parameters increase the total number of FLOPs in some cases, they offer markedly higher accuracy, leading to a favourable trade-off between computational cost and classification performance. Notably, Tempotron-based models achieve the highest throughput, whereas BP- and STDP-based LIF-SNNs provide the best accuracy-to-compute ratio.
In addition to overall accuracy, Tables 2 and 4 include class-wise precision, recall, and F1-scores, which provide a more detailed picture of model behaviour. The extended configurations, in which intrinsic neuron parameters are optimized jointly with synaptic weights, consistently lead to more balanced class-wise performance. This is reflected in higher and more symmetric F1-scores across classes, indicating that the models are not only more accurate overall but also less biased toward a particular class. In contrast, baseline configurations optimized only over network size and learning rate often exhibit strong class imbalance, with one class being detected almost perfectly (recall close to $\approx$ 1.0) while the other collapses to near-zero precision or recall. This behaviour is particularly evident for STDP and Tempotron baselines, where the model tends to favour the more frequent or more separable class. Learning intrinsic parameters alleviates this issue by enabling the neurons to adjust their internal timescales and thresholds, thereby improving separability in both classes. Overall, the class-wise metrics confirm that the accuracy improvements observed in the extended configurations stem not only from higher correct classification rates but from a more robust and balanced decision boundary across classes.
@@
Backpropagation and STDP benefited the most from intrinsic parameter optimization, yielding accuracy gains of up to 13.50 percentage points and 10.00 percentage points over their respective baselines in the LIF-SNN, and up to 8.50 and 7.00 percentage points in the meta-SNN. Tempotron, while already showing a strong baseline performance (96.00% for LIF-SNN and 89.50% for meta-SNN), exhibited smaller improvements of about 3.50 percentage points. This suggests that its performance relies primarily on synaptic weight dynamics rather than intrinsic parameter tuning. As a result, parameter optimization (thresholds and decay constants) contributes less to performance for Tempotron compared to Backpropagation or STDP. Thus, the optimization configurations and their interpretation are summarized in Table 6.